# Diorama: Unleashing Zero-shot Single-view 3D Scene Modeling

### v1

@@ 4 Experiments
Figure 6: Examples on SSDB images. Scenes are modelled from estimated metric depth by Metric3Dv2 [33] and scene graphs are generated using GPT-4o. The third column visualizes the support hierarchy of the generated scene graphs. We predict object placement and arrangement in 3D. The “Arrangement 1” column uses ground-truth 3D shapes. “Arrangement 2” and “Arrangement 3” present similar semantic arrangements using different retrieved 3D shapes.
Ground Truth Depth Estimated Depth
Zero-shot pose estimation Scene-aware Alignment $\uparrow$ Collision $\downarrow$ Relation $\uparrow$ Scene-aware Alignment $\uparrow$ Collision $\downarrow$ Relation $\uparrow$
rAcc tAcc sAcc Acc rAcc tAcc sAcc Acc
Best-matching multiview 0.34 0.93 0.52 0.19 11.43 0.58 0.34 0.84 0.44 0.15 12.29 0.25
ZSP [30] w/ DinoV2 0.37 0.93 0.57 0.23 10.48 0.59 0.42 0.85 0.55 0.26 8.29 0.25
ZSP [30] w/ ft DinoV2 0.45 0.94 0.66 0.34 7.90 0.61 0.41 0.85 0.54 0.24 8.46 0.24
GigaPose [63] 0.36 0.95 0.71 0.27 7.91 0.61 0.36 0.84 0.64 0.22 7.44 0.22
Ours 0.45 0.95 0.71 0.33 7.26 0.61 0.41 0.84 0.64 0.25 7.37 0.25
Table 1: Comparison of zero-shot pose estimation methods for the 9-DoF CAD alignment task on SSDB images. GT objects are used to avoid 2D perception error for system analysis. We present results given GT and estimated depth.
Ablation Collision $\downarrow$ Scene Structure $\uparrow$
orientation placement overall
w/o optimization 7.26 0.18 0.01 0.00
+ Orientation 6.34 0.97 0.01 0.01
+ O, Placement 7.61 0.97 0.94 0.91
+ O, P, Space 5.90 0.97 0.54 0.53
+ O, P, S, Refinement 4.43 0.97 0.95 0.92
w/ GT scene graph 3.28 0.97 0.97 0.95
w/ GT architecture 4.15 0.97 0.97 0.95
@@ 4.1 Experimental setup
Datasets. We evaluate on the Stanford Scene Database (SSDB) [23], which contains around 130 scenes. We choose this dataset as the scenes represent highly cluttered arrangements such as office desk setups and chemistry laboratories annotated using 1,700 unique CAD models ranging from large furniture (e.g. chair, table, bookshelf) to small items (e.g. handbag, clothing hanger, doll). We use Blender to render 344 scene images, with 3 camera views per scene on average. The SSDB dataset provides a rich support hierarchy annotation and defines the room architecture, allowing us to evaluate these two aspects in our output. To enlarge the 3D shape repository beyond what is present in SSDB, we include shapes from HSSD [44] and the LVIS set of Objaverse [14] to form an out-of-distribution (OOD) collection of 60K 3D shapes for retrieval.
Metrics. We evaluate the components of Diorama with a suite of metrics for object alignment, overall scene quality, and 3D shape retrieval. For alignment, we report the alignment accuracy for translation (tAcc), rotation (rAcc), and scale (sAcc), as well as overall accuracy (Acc). Prior work [31, 48, 27] judged a CAD model as correctly aligned if the translation error $\leq 20$cm, the geodesic rotation error $\leq 20^{\circ}$, and the scale ratio $\leq 20\%$. However, we found that scene plausibility can be severely reduced due to a few objects with unreasonable position, orientation or size even when average accuracy is high. Thus, we introduce a scene-aware alignment accuracy to penalize such cases. An object is considered correctly aligned under scene awareness if its relative position deviation $\leq 20^{\circ}$, relative rotation error $\leq 20^{\circ}$ and relative scale error $\leq 0.2$ to every other object in the scene. To measure the overall scene quality, we decompose the evaluation into two aspects: average mesh collision and scene structure hierarchy where we calculate the support relation prediction accuracy (where a prediction is correct if both support object and surface assignment are correct) and supportness accuracy where whether objects are properly positioned. To measure 3D shape retrieval similarity against the ground-truth shape, we compute the chamfer distance (CD) as in prior work [90, 27].
4.2 Quantitative Evaluation on SSDB
@@
Table 3: Retrieval similarity (CD) results on SSDB images with top-5 retrievals, multiplied by $10^{3}$ for readability. We compare retrieving objects with different models - CLIP, OpenShape (OS), DuoduoCLIP (DD) - from SSDB (SS) and a wider set of out-of-distribution objects (OOD). We take the candidate with the minimum L1 Chamfer Distance to the ground truth. “T” and “V” mean using text or visual modality for query only respectively. “H” stands for using both modalities.
@@
Figure 7: Qualitative results on ScanNet images. As detection, depth estimation and shape retrieval are modular components in both DiffCAD and Diorama, we use the same mask, depth and 3D shape inputs for alignment for fair comparison. We handle more categories in 1st and 4th examples (laptop and backpack).
Input Output Input Output
@@
Figure 8: Examples for in-the-wild images. The output scenes capture large and small objects, and represent complex relations such as objects on walls (e.g., wall clock in top left) and objects on other objects (vase with flower on table, in the top right).
Text prompt Input Output Text prompt Input Output
@@
Figure 9: Examples of applying Diorama in a text-to-scene setting. We generate an input image from a text prompt using Flux-1, and apply our system to the image. The output scenes conform to the text prompt.
@@
We present evaluations of retrieval, object pose, and scene optimization. See the supplement for evaluation of the architecture reconstruction, and more ablations.
@@
9D CAD alignment to image object. We compare our zero-shot pose estimation module with representative methods, including ZSP [30] and GigaPose [63], on the 9-DoF CAD alignment task on SSDB images. Since GigaPose originally cannot predict the size of CAD model, we augment it by integrating depth information. We include a baseline which only performs coarse pose selection using the best-matching multiview image with the highest average correspondence similarity. For ZSP, we use DinoV2 as the default feature extractor rather than Dino [6] for fair comparison. We also experiment with a variant of ZSP leveraging a fine-tuned DinoV2. Tab. 1 shows that our methods perform better with both ground-truth and estimated depth, as measured by metrics for alignment, collision and support relation prediction. ZSP performs slightly better on the overall alignment metric but significantly falls behind on the scale metric, which leads to implausible scene layouts in the end. The failure on scale estimation is due to the RANSAC-based pose solver that is substantially affected by noisy point correspondences.
@@
3D shape retrieval results. Tab. 3 evaluates our multimodal 3D shape retrieval approach against state-of-the-art for open-world methods, including CLIP [69] and OpenShape [57]. We use the ViT-bigG-14 checkpoint of CLIP. We present results both for retrieving from a relatively small set of SSDB shapes, and from a large set of OOD shapes. Our approach outperforms others by a big margin reflecting that multiple modalities compensate for potential ambiguity in language or visual information loss in a photo of a complicated scene. The inferior performance of OpenShape using the same hybrid retrieval strategy implies that the multiview representation of 3D shapes we use is a better fit for the open-world problem setting.
@@ 4.3 Zero-shot Results on ScanNet
We validate our zero-shot open-world system on real-world images from ScanNet25k [12] and ShapeNet [7] CAD model annotations provided by Scan2CAD [4]. We deliberately evaluate on cluttered environments capturing occlusion and multiple real world objects. We do this by selecting 600 images from the standard evaluation set where each image contains at least 3 different object categories or 5 object instances. These images contain around 3,000 instances spanning 24 categories, compared to the evaluation set in prior work [42] with 2,500 instances from 2,100 images covering 6 categories. We compare to the state-of-the-art prior work ROCA [31] and DiffCAD [27] where ROCA is a fully-supervised deterministic end-to-end model trained on training data of Scan2CAD and DiffCAD is a weakly-supervised probabilistic model trained on generated synthetic data. However, we note that both these methods cannot scale up and generalize to unobserved real-world images as ROCA consumes expensive real-world data for training and DiffCAD requires training for each category individually and inefficient inference due to a denoising network design. In contrast, Diorama provides a training-free solution showing comparable performance against DiffCAD and more flexible capability against ROCA. We show a qualitative comparison of CAD alignment in Fig. 7. Note that we can handle out-of-distribution objects (laptop and backpack) in the first and fourth row, which ROCA and DiffCAD cannot process due to limited training data. See the supplement for quantitative results.
4.4 In-the-wild Results
Real-world image. We select a set of real-world images representing real-world object arrangement from pixabay, Pexels and Unsplash. We show qualitative results in Fig. 8.
Text-to-scene. We further expand the capability of Diorama to the text-to-scene problem. We first generate high-resolution realistic image given a text prompt using Flux-1 and apply our system to the image. We show qualitative results in Fig. 9.
@@ 5 Conclusion
We introduced Diorama, the first zero-shot open-world system that can holistically model a 3D scene from an RGB image without requiring end-to-end training or human annotations. Our system is composed of robust, generalizable solutions to each subtask of our problem: architecture reconstruction, 3D shape retrieval, 9D object pose estimation, and scene layout optimization. We evaluate our system on synthetic images and real-world data with quantitative analysis and visual comparisons, and demonstrate generalization to internet images and text-to-scene applications.
Acknowledgements. This work was funded in part by a CIFAR AI Chair, a Canada Research Chair, NSERC Discovery Grant, and NSF award #2016532. Daniel Ritchie is an advisor to Geopipe and owns equity in the company. Geopipe is a start-up that is developing 3D technology to build immersive virtual copies of the real world with applications in various fields, including games and architecture. We thank Xiaohao Sun, Jiayi Liu, Xingguang Yan, Chirag Vashist, and Wuyang Chen for feedback on the paper draft.
@@ A.2 LMM-powered scene graph generation
After obtaining the set of objects in the scene, we then use an LMM (e.g. GPT-4o) to obtain the scene-graph. The scene-graph provide information about relationships between objects, including support relations.
We specify to the LMM the desired output response in JSON format so it can be successfully executed by the program afterward. We also limit object supporting relations to be selected from [“placed on”, “mounted on”].
To encourage the LMM to obey the provided numeric object marks and reduce hallucination errors, we use interleaved text prompt by incorporating the object marks and template response format into the scene graph generation prompt directly for symbolic reference.
@@ A.3 Architecture reconstruction
To obtain a planar reconstruction of the architecture, we apply a three step process: 1) segmenting out objects and inpainting to obtain empty rooms, 2) using depth estimation to obtain 3D points, and 3) clustering points by normals and plane fitting to obtain planar architecture surfaces.
@@
Once we have obtained the point clusters, we fit bounded planar segments to the points to obtain the final architecture. We start by fitting a plane using RANSAC  [22]to obtain a plane equation, and then estimate the bounding boxes of architecture elements. To obtain tight boxes, we compute convex hulls of point cloud segments, project the convex hull vertices onto 2D plane and apply rotating calipers algorithm  [73] to obtain the bounding box with the smallest area. Finally, we convert vertices back into 3D. While such initialization provides good plane bounds, we need to refine the plane further to account for imperfections of previous steps. We begin by making sure that all the walls are orthogonal to the floor, this is simply done by solving a system of linear equations that finds a vector orthogonal to the intersection line of floor and wall and lies on the floor plane. We proceed to use this vector as a new normal for our wall equation. Then we proceed to refine the plane bounds by finding the intersections of each architecture element, projecting two closest vertices of one plane onto the intersection line and adjusting the vertices of the plane we projected onto to the projected vertices of the other plane, effectively connecting them. Finally, we duplicate the vertices with a small offset along the direction opposite to normal and export the meshes of architectural components.
Figure 12: Qualitative results on architecture reconstruction.
RaC [75] DAv2 236 40.3 (58.8) 39.8 (12.2) 23.2 0.67 0.645 43
RaC [75] M3D 258 45.2 (60.3) 33.7 (11.6) 21.0 1.23 0.908 49
PlainRecon DAv2 344 47.8 13.3 27.5 1.26 0.503 23
PlainRecon M3D 344 58.6 9.6 18.9 1.37 0.447 29
Table 4: General architecture plane reconstruction results on SSDB. * and indicate estimated depth from Metric3DV2 [33] and DepthAnythingV2 [95] respectively. Numbers in parentheses indicate results of RaC when only consider successful predictions.
@@
Household 466 / 5577 0.44 0.84 0.71 0.29
Furniture 24 / 1137 0.53 0.84 0.69 0.34
Occluded 369 / 3175 0.43 0.82 0.64 0.25
Complete 401 / 3539 0.48 0.86 0.77 0.34
Supported 468 / 5715 0.46 0.85 0.71 0.30
Supporting 79 / 1634 0.46 0.82 0.70 0.28
@@
# retrievals Ground Truth Depth Estimated Depth
rAcc tAcc sAcc Acc rAcc tAcc sAcc Acc
1 0.21 0.93 0.46 0.09 0.21 0.82 0.42 0.08
4 0.32 0.96 0.64 0.18 0.32 0.85 0.59 0.15
8 0.37 0.96 0.69 0.21 0.36 0.84 0.65 0.18
Table 7: 9D CAD alignment results give a different number of retrieved 3D shapes using either ground-truth or estimated depth.
Ablation Collision $\downarrow$ Scene Structure $\uparrow$
orientation placement overall
all-in-one 12.49 0.18 0.84 0.15
stage-wise 4.43 0.97 0.95 0.92
@@
Figure 14: More examples on SSDB images.
@@
Figure 15: More examples for in-the-wild images.
Text prompt Input Output Text prompt Input Output

### v2

@@ 4 Experiments
Figure 6: Comparison examples on SSDB. We capture architecture details, object geometry matching and spatial relations better.
Figure 7: Scene modeling examples of Diorama on SSDB images. The third column visualizes the support hierarchy of the generated scene graphs. We predict object placement and arrangement in 3D. The “Arrangement 1” column uses ground-truth 3D shapes. “Arrangement 2/3” presents similar semantic arrangements using different retrieved 3D shapes.
Method Scene-aware Alignment $\uparrow$ CD $\downarrow$ User $\uparrow$ System Cost $\downarrow$
rAcc tAcc sAcc Acc API (\$) Time (min)
ACDC [11] 0.20 0.56 0.36 0.04 14.0 14.9 1.44 23.2
Ours 0.23 0.68 0.49 0.08 9.5 85.1 0.12 3.7
Table 1: System comparison on SSDB images. We exclude GT 3D shapes from retrieval for fair comparison. “User” indicates the user preference over generated scenes by two methods.
@@ 4.1 Experimental setup
Datasets. We demonstrate on the Stanford Scene Database (SSDB) [20], which contains around 130 scenes. It provides a rich support hierarchy annotation and defines the room architecture, allowing us to evaluate our output comprehensively. The SSDB scenes represent highly cluttered arrangements such as office desk setups and chemistry laboratories annotated using 1,722 unique CAD models ranging from large furniture (e.g. chair, table, bookshelf) to small items (e.g. handbag, clothing hanger, doll). We use Blender to render 344 scene images, with 3 camera views per scene on average. To enlarge the 3D shape repository, we include shapes from HSSD [38] and the LVIS set of Objaverse [12] (Obja-LVIS) to form an out-of-distribution (OOD) collection of 60K 3D shapes for retrieval.
Metrics. We evaluate the different aspects of Diorama with a suite of metrics. For architecture reconstruction, we report 2D IoU, Pixel Error (PE), Edge Error (EE) and Root Mean Squared Error (RMSE) following [68]. We also introduce bounding-box-centered chamfer distance (CDb) to measure the correctness of 3D structure of the predicted architecture. To evaluate object arrangement, we introduce a scene-aware alignment accuracy (Acc) to eliminate the effect of inaccurate depth prediction and penalize cases where scene plausibility is severely reduced due to a few objects with unreasonable poses. An object is considered correctly aligned under scene awareness if its relative translation deviation $\leq 20^{\circ}$ (tAcc), relative rotation error $\leq 20^{\circ}$ (rAcc) and relative scale error $\leq 0.2$ (sAcc) to every other object in the scene. To measure the overall scene quality, we decompose the evaluation into two aspects: average mesh collision and scene structure hierarchy where we calculate the support relation prediction accuracy (where a prediction is correct if both support object and surface assignment are correct) and supportness accuracy where whether objects are properly positioned. To measure 3D shape retrieval similarity against the ground-truth shape, we compute the chamfer distance (CD) as in prior work [81, 24], multiplied by $10^{3}$ for readability.
4.2 Zero-shot Results on SSDB
Figure 8: Examples of planar architecture reconstruction on SSDB and real-world images.
RaC [68] DAv2 236 40.3 (58.8) 39.8 (12.2) 23.2 0.67 0.645 43
RaC M3Dv2 258 45.2 (60.3) 33.7 (11.6) 21.0 1.23 0.908 49
ACDC [11] DAv2 344 46.8 17.0 31.0 1.54 0.563 32
PlainRecon DAv2 344 47.8 13.3 27.5 1.26 0.503 23
PlainRecon M3Dv2 344 58.6 9.6 18.9 1.37 0.447 29
Table 2: Planar architecture reconstruction results on SSDB. Numbers in parentheses indicate results of RaC when only considering successful predictions.
@@ 4.2 Zero-shot Results on SSDB
Table 3: Retrieval similarity (CD) results on SSDB images. We compare different models - CLIP, OpenShape (OS), DuoduoCLIP (DD) - retrieving from SSDB (SS) and out-of-distribution (OOD) objects. “T” and “V” mean using text or visual modality for query only respectively. “H” stands for using both modalities.
Method Scene-aware Alignment $\uparrow$ Collision $\downarrow$ Relation $\uparrow$
rAcc tAcc sAcc Acc
BM baseline 0.34 0.93 0.52 0.19 11.43 0.58
ZSP [26] 0.36 0.92 0.59 0.25 8.76 0.60
ZSP w/ DinoV2 0.37 0.85 0.56 0.26 9.59 0.59
ZSP w/ ft DinoV2 0.38 0.93 0.58 0.25 10.58 0.44
GigaPose [56] 0.36 0.95 0.71 0.27 7.91 0.61
Ours 0.47 0.95 0.70 0.37 6.42 0.62
Table 4: Comparison of zero-shot pose estimation methods for the 9-DoF CAD alignment task on SSDB images. GT objects and depth are used to avoid 2D perception error for system analysis.
Ablation Collision $\downarrow$ Scene Structure $\uparrow$
orientation placement overall
w/o optimization 6.42 0.19 0.01 0.00
+ Orientation 5.17 0.98 0.01 0.01
+ O, Placement 8.08 0.98 0.93 0.91
+ O, P, Space 6.08 0.98 0.54 0.54
+ O, P, S, Refinement 3.78 0.98 0.95 0.93
w/ GT scene graph 3.12 0.98 0.94 0.92
w/ GT architecture 3.88 0.96 0.93 0.90
@@
Figure 9: Examples on ScanNet images. We use the same mask, depth and 3D shape inputs for DiffCAD and ours for fair comparison as they are both modular methods. We handle more categories in 1st and 4th examples (laptop and backpack).
Input Output Input Output Input Output
@@
Figure 10: Examples on real-world internet images. The output scenes capture large and small objects, and represent complex relations such as objects on walls (e.g., wall clock in top left) and objects on other objects (vase with flower on table, in the top right).
Text prompt Generated image Output
@@
Figure 11: Examples of applying Diorama in a text-to-scene setting. The output scenes conform to the text prompt.
@@
We present evaluations of overall system performance, architecture reconstruction, shape retrieval, object pose estimation and scene optimization. See the supplement for more ablations. Fig. 7 shows multiple semantic-similar arrangements produced by Diorama on scenes of complex layout hierarchy.
System comparison. We compare with ACDC [11], another modular system of building synthetic scenes from images for robot policy learning. We evaluate the overall scene modeling performance over metrics of object alignment, similarity and system cost. We exclude ground-truth SSDB 3D shapes from retrieval for fair comparison. Tab. 1 shows that Diorama outperforms ACDC noticeably. Both methods present relatively low performance on alignment accuracy due to object retrieval mismatch and undesired depth prediction. While both approaches leverage LLM, ours is ×$\times 12$ cheaper and ×$\times 6$ faster than ACDC. We also conduct a user study in terms of object matching and scene quality on sampled reconstructed scenes. We achieve 85.1% user preference on outputs generated by our method. We show qualitative examples in Fig. 6 and note that Diorama reconstructs more complete and fine-grained room structure, and captures better object arrangements. Since ACDC heavily relies on a physics engine to solve the collision and floating problem, it cannot explicitly place an object on certain supporting surface. See more examples in the supplement.
Architecture plane reconstruction. We compare our proposed PlainRecon with RaC and ACDC in Tab. 2, with predicted depth from DepthAnythingV2 (DAv2) and Metric3DV2 (M3Dv2). We find that RaC fails in roughly 30% of the cases due to infeasible solutions produced by the constrained discrete optimization procedure, while PlainRecon and ACDC are able to handle all inputs. PlainRecon remarkably outperforms RaC and ACDC on most of the metrics, while RMSE being an outlier where PlainRecon underperforms significantly. The superior performance of PlainRecon on CDb means that it captures the 3D room structure of the scene better than others. Our results also show that M3Dv2 depth is overall superior to DAv2 for architecture reconstruction. We visualize both reconstruction results on SSDB and internet images in Fig. 8.
3D shape retrieval results. Tab. 3 evaluates our multimodal 3D shape retrieval approach against CLIP [62] and OpenShape [50]. We use the ViT-bigG-14 checkpoint of CLIP. We present results both for retrieving from SSDB shapes, and from a larger set of OOD shapes. We retrieve top-5 objects and take the candidate with the minimum L1 Chamfer Distance to the ground truth. Our approach outperforms others by a big margin reflecting that multiple modalities compensate for potential ambiguity in language or visual information loss in a photo of a complicated scene. The inferior performance of OpenShape implies that the multiview representation of 3D shapes we use is a better fit for the open-world problem setting.
@@
9D CAD alignment to image object. We compare our zero-shot pose estimation module with representative methods, including ZSP [26] and GigaPose [56], on the 9-DoF CAD alignment task on SSDB images. Since GigaPose originally cannot predict the size of CAD model, we augment it by integrating depth information. We include a baseline (BM) which only performs coarse pose selection using the best-matching multiview image with the highest average correspondence similarity. For ZSP, we experiment with variants in terms of different feature extractors, including the default Dino [5], DinoV2 and a fine-tuned DinoV2 backbone from GigaPose. Tab. 4 shows that our method performs the best as measured by metrics for overall alignment, collision and support relation prediction, while achieving comparable performance on sAcc against GigaPose. The failure on scale estimation is due to the RANSAC-based pose solver that is substantially affected by noisy point correspondences. See additional results on estimated depth in the supplement.
@@ 4.3 Zero-shot Results on ScanNet
We validate our zero-shot open-world system on real-world images from ScanNet25k [10] and ShapeNet [6] CAD model annotations provided by Scan2CAD [4]. We deliberately evaluate on cluttered environments capturing occlusion and multiple objects. We select 600 images from the standard evaluation set where each image contains at least 3 different object categories or 5 object instances. These images contain around 3,000 instances spanning 24 categories, compared to the evaluation set in prior work [27] with 2,500 instances from 2,100 images covering 6 categories. We compare to ROCA [27] and DiffCAD [24] where ROCA is a fully-supervised deterministic end-to-end model trained solely on Scan2CAD and DiffCAD is a weakly-supervised probabilistic model trained on generated synthetic data. Both prior works cannot scale up and generalize to unobserved real-world images due to expensive data acquisition, infeasibility of training individual models for every category or inefficient inference. In contrast, Diorama provides a training-free solution showing comparable performance against DiffCAD and more flexible capability against ROCA. We show a qualitative comparison of CAD alignment in Fig. 9. Note that we can handle out-of-distribution objects (laptop and backpack) in the first and fourth row, which ROCA and DiffCAD cannot process due to limited data. See the supplement for more results.
4.4 Zero-shot Results on in-the-wild Images
Real-world internet images. We select a set of real-world internet images representing real-world object arrangement from pixabay, Pexels and Unsplash. See Fig. 10 and more examples in the supplement.
Text-conditioned generated images. We further expand the capability of Diorama to the text-to-scene problem. We first generate high-resolution realistic image given a text prompt using Flux-1 and apply our system to the image. See Fig. 11 and more examples in the supplement.
@@ 5 Conclusion
We introduced Diorama, the first zero-shot open-world system that can holistically model a 3D scene from an RGB image without requiring end-to-end training or human annotations. Our system is composed of robust, generalizable solutions to each subtask of our problem: architecture reconstruction, 3D shape retrieval, 9D object pose estimation, and scene layout optimization. We evaluate our system on synthetic images and real-world data with quantitative analysis and visual comparisons, and demonstrate generalization to real-world images and text-to-scene applications.
Acknowledgements. This work was funded in part by a CIFAR AI Chair, a Canada Research Chair, NSERC Discovery Grant, and NSF award #2016532. Daniel Ritchie is an advisor to Geopipe and owns equity in the company. Geopipe is a start-up that is developing 3D technology to build immersive virtual copies of the real world with applications in various fields, including games and architecture. We thank Xiaohao Sun, Jiayi Liu, Xingguang Yan, Chirag Vashist, Wuyang Chen, Dongchen Yang and Mahdi Miangoleh for feedback on the paper draft.
@@ A.2 LMM-powered scene graph generation
[…]
We specify to the LMM the desired output response in JSON format so it can be successfully executed by the program afterward. We also limit object supporting relations to be selected from [“placed on”, “mounted on”]. To encourage the LMM to obey the provided numeric object marks and reduce hallucination errors, we use interleaved text prompt by incorporating the object marks and template response format into the scene graph generation prompt directly for symbolic reference.
@@ A.3 Architecture reconstruction
To obtain a planar reconstruction of the architecture, we apply a three-step process: 1) segmenting out objects and inpainting to obtain empty rooms, 2) using depth estimation to obtain 3D points, and 3) clustering points by normals and plane fitting to obtain planar architecture surfaces.
@@
Once we have obtained the point clusters, we fit bounded planar segments to the points to obtain the final architecture. We start by fitting a plane using RANSAC [19] to obtain a plane equation, and then estimate the bounding boxes of architecture elements. To obtain tight boxes, we compute convex hulls of point cloud segments, project the convex hull vertices onto 2D plane and apply the rotating calipers algorithm  [66] to obtain the bounding box with the smallest area. Finally, we convert vertices back into 3D. While such initialization provides good plane bounds, we need to refine the plane further to account for imperfections of previous steps. We begin by making sure that all the walls are orthogonal to the floor, this is simply done by solving a system of linear equations that finds a vector orthogonal to the intersection line of floor and wall and lies on the floor plane. We proceed to use this vector as a new normal for our wall equation. Then we proceed to refine the plane bounds by finding the intersections of each architecture element, projecting two closest vertices of one plane onto the intersection line and adjusting the vertices of the plane we projected onto to the projected vertices of the other plane, effectively connecting them. Finally, we duplicate the vertices with a small offset along the direction opposite to normal and export the meshes of architectural components.
Method Scene-aware Alignment $\uparrow$ Collision $\downarrow$ Relation $\uparrow$
rAcc tAcc sAcc Acc
BM baseline 0.34 0.84 0.44 0.15 12.29 0.58
ZSP [26] 0.32 0.88 0.50 0.18 8.87 0.59
ZSP w/ DinoV2 0.33 0.81 0.48 0.20 9.80 0.57
ZSP w/ ft DinoV2 0.34 0.83 0.50 0.21 9.97 0.59
GigaPose [56] 0.36 0.89 0.67 0.24 7.43 0.61
Ours 0.41 0.91 0.61 0.28 6.93 0.61
Table 6: Additional comparison of zero-shot pose estimation methods for the 9-DoF CAD alignment task on SSDB images using estimated depth by Metric3DV2.
@@
Household 466 / 5577 0.46 0.87 0.71 0.32
Furniture 24 / 1137 0.54 0.86 0.71 0.39
Occluded 369 / 3175 0.44 0.84 0.62 0.28
Complete 401 / 3539 0.50 0.89 0.78 0.38
Supported 468 / 5715 0.48 0.87 0.71 0.34
Supporting 79 / 1634 0.46 0.84 0.69 0.31
@@
# retrievals Ground Truth Depth Estimated Depth
rAcc tAcc sAcc Acc rAcc tAcc sAcc Acc
1 0.22 0.92 0.52 0.11 0.21 0.88 0.47 0.10
4 0.32 0.96 0.68 0.20 0.32 0.93 0.63 0.18
8 0.37 0.97 0.72 0.23 0.37 0.94 0.68 0.21
Table 8: 9D CAD alignment results given a different number of retrieved 3D shapes using either ground-truth or estimated depth.
Ablation Collision $\downarrow$ Scene Structure $\uparrow$
orientation placement overall
all-in-one 12.80 0.19 0.92 0.16
stage-wise 3.78 0.98 0.95 0.93
@@
Figure 15: More comparison examples on SSDB.
Figure 16: More Diorama examples on SSDB images.
@@
Figure 17: More examples for real-world internet images.
Text prompt Generated image Output Text prompt Generated image Output